Which equation will result if the substitution method is used to solve the following system of linear equations?
x + y = 6
x = y + 2

Answer by LinnW(1048)

replace x with (y+2) in the equation x + y = 6
(y+2) + y = 6
2y + 2 = 6
add -2 to each side
2y = 4
y = 2
Now substitute 2 for y in x = y + 2
x = 2 + 2
x = 4